# High-fidelity Metalens Imaging via Aberration Correcting Transformers

### v1

@@ Abstract
Metalens is an emerging optical system with an irreplaceable merit in that it can be manufactured in ultra-thin and compact sizes, which shows great promise of various applications such as medical imaging and augmented/virtual reality (AR/VR). Despite its advantage in miniaturization, its practicality is constrained by severe aberrations and distortions, which significantly degrade the image quality. Several previous arts have attempted to address different types of aberrations, yet most of them are mainly designed for the traditional bulky lens and not convincing enough to remedy harsh aberrations of the metalens. While there have existed aberration correction methods specifically for metalens, they still fall short of restoration quality. In this work, we propose MetaFormer, an aberration correction framework for metalens-captured images, harnessing Vision Transformers (ViT) that has shown remarkable restoration performance in diverse image restoration tasks. Specifically, we devise a Multiple Adaptive Filters Guidance (MAFG), where multiple Wiener filters enrich the degraded input images with various noise-detail balances, enhancing output restoration quality. In addition, we introduce a Spatial and Transposed self-Attention Fusion (STAF) module, which aggregates features from spatial self-attention and transposed self-attention modules to further ameliorate aberration correction. We conduct extensive experiments, including correcting aberrated images and videos, and clean 3D reconstruction from the degraded images. The proposed method outperforms the previous arts by a significant margin. We further fabricate a metalens and verify the practicality of MetaFormer by restoring the images captured with the manufactured metalens in the wild. Code and pre-trained models are available at https://benhenryl.github.io/MetaFormer.
@@ 1 Introduction
Metalens is an advanced optical device that leverages nanoscale engineering to manipulate light. Unlike traditional lenses relying on glass or plastic surfaces, the metalens harnesses a metasurface, a thin optical element consisting of nanostructures modulating optical properties such as phase, amplitude, and polarization. It can be fabricated in ultra-thin and compact sizes compared to the traditional bulky lens, making it an ideal option for various domains, including the medical industries, consumer electronics, and augmented/virtual reality (AR/VR) [95, 75, 87, 31, 67, 54, 43, 52]. Despite its great potential for miniaturization, existing metalenses have faced critical challenges, such as aberrations and distortions in the captured images, hampering their broad uses across diverse applications.
Aberrations in optical systems occur due to practical constraints and physical properties of lens materials. Different types of aberrations arise for different reasons, e.g. chromatic aberrations are produced when lights with different wavelengths refract at slightly different angles when they pass through a lens, resulting in being focused at different points. Another type of aberrations, spherical aberrations, occur when rays transmitting through the edges and center of the lens focus at different locations.
@@
Several previous works [10, 23, 28] have attempted to correct these aberrations, primarily addressing them for general commercial cameras and not effective enough to mitigate intense aberrations of metalens. On the other hand, Tseng et al. [80] and Chakravarthula et al. [9] devised aberration correcting methods for the metalens. Tseng et al. [80] adopted convolutional neural networks (CNNs) and Chakravarthula et al. [9] employed a diffusion model [36, 74] for restoration. While promising, these methods have yet to reach the desired restoration quality.
Besides the aberration corrections, numerous studies have tackled other image degradations such as blur, noise, raindrops, haze, etc. A line of research introduced [39, 65, 46, 47] generative models [29, 74, 36] to produce clean images from degraded images, and other arts [60, 93, 63, 32, 49, 91] adopted CNNs. Another line of works [92, 55, 13, 16, 81] harnessed Vision Transformer (ViT) [22] for image restoration. Although they have shown prominent results, straightforwardly applying them to metalens-induced aberration correction is undesirable. This is because metalens can manipulate multiple optical properties, resulting in complex aberration patterns that vary in response to different wavelengths, which lead to severe chromatic aberrations. Moreover, aberrations in metalens images can amplify noise and distort intensity variations across channels. These aberrations are more complicated than the degradations encountered in conventional image restoration tasks, and those approaches devised for general image restoration lack in capacity to adapt to these distortions.
Among diverse frameworks, ViT-based approaches have proven effective for low-level computer vision tasks. ViT can effectively capture global context by leveraging a self-attention that allows each pixel in an image to communicate with all others, unlike CNNs, where pixels can interact only with neighboring pixels. This capability for modeling global relationships is especially crucial for image restoration, where understanding the overall structure plays a key role in removing noise or artifacts while preserving intricate details. Moreover, self-attention modules enable dynamic weighting of features, empowering the model to adaptively focus on the most relevant regions. While ViT has demonstrated its competence in various tasks, its potential for restoring metalens images remains largely unexplored.
It motivated us to propose MetaFormer, an aberration correction Transformer tailored for correcting the aberration of metalens-captured images. Particularly, we design a Multiple Adaptive Filters Guidance (MAFG) that employs several Wiener filters [86], which have been widely used to reduce noise in signals. The advantages of MAFG are threefold: first, the representations deconvolved with different Wiener filters show diverse degrees of noise suppression and recovered fine details, allowing them to be processed complementarily in the restoration network. Secondly, MAFG provides a mechanism for handling unknown noise distributions. The Wiener filter penalizes noise based on the signal-to-noise ratio (SNR), which is usually obscure in real-world scenarios. By incorporating multiple Wiener filters with varied noise suppression parameters, MAFG can effectively address unknown noise distributions. Lastly, we adapt these parameters based on the image intensity, allowing for flexible handling across a broad range of images.
Furthermore, we extend Restormer [92] by incorporating both spatial and transposed self-attention and aggregating features from different attention modules, proposing a Spatial and Transposed self-Attention Fusion (STAF) module. Various previous arts [13, 16, 81] adopted encoder-decoder architecture with two types of attention mechanisms (spatial and transposed), but they apply them alternatively in both encoder and decoder, though they have different roles. The encoder focuses on capturing global structures, whereas the decoder emphasizes more on reconstructing fine details of images. Motivated by this, we implement spatial and transposed self-attention in parallel and fuse their features to guide the encoder in capturing global context and the decoder in enhancing fine reconstruction.
@@
We conducted extensive experiments across various imaging scenarios to validate the effectiveness of the proposed MetaFormer. First, on synthetic aberrated images generated using point spread functions (PSF) from real metalenses, MetaFormer demonstrated notable performance improvements, outperforming the baseline model by 3.24 dB in PSNR and achieving state-of-the-art restoration quality across multiple evaluation metrics. Second, we extended and applied the proposed techniques to a video restoration model, where results further confirmed the proposed technique’s superiority in aberration correction in videos. Third, we tested MetaFormer on 3D reconstruction tasks using multi-view images (3D Gaussian Splatting [41]), demonstrating its capability to maintain both high-quality aberration correction and 3D consistency across views. Finally, we fabricated metalens using the open-source point spread function (PSF) of a metalens [80] and evaluated the proposed method on the images captured with the manufactured metalens. The corrected images demonstrated promising results, revealing new possibilities in metalens imaging.
To sum up, our contributions are as follows:
@@
  • We enhanced the baseline model by 3.24 dB in PSNR, accomplishing state-of-the-art restoration performance.
@@ 2.1 Metalens
2.2 Image Restoration models
Image restoration aims to recover high-quality images from degraded observations. Traditionally, Wiener deconvolution [86] recovered images by solving complex inverse problems. However, it often struggled with noise and complex degradations, limiting their real-world applications. Afterward, deep learning has revolutionized image restoration with data-driven methods. Vision Transformer, in particular, (ViT) [22] has achieved remarkable performances in image restoration tasks, including denoising [81, 11, 85, 26, 92, 53], deblurring [85, 92, 59, 45, 79], and super resolution [14, 15, 55, 77, 61, 16], by exploiting self-attention that can capture global interactions effectively.
Restormer [92] stood out with its superior performance, introducing transposed self-attention mechanisms that apply self-attention across the channel of an image, enabling efficient processing of high-resolution images. However, it might limit the exploitation of spatial information. X-Restormer [13] further incorporated spatial self-attention mechanisms into Restormer to handle these shortcomings. This allowed the model to better capture spatial dependencies, still there is room for improvement in that it deployed attention blocks identically for encoder and decoder though they have different roles when learning the representation.
Figure 2: The overview of MetaFormer. It comprises Multiple Adaptive Filters Guidance (MAFG) which produces different representations with various noise-detail balances, and a Spatial and Transposed self-Attention Fusion (STAF) module that aggregates features differently in encoder and decoder.
2.3 Aberration correction models
@@
Blind aberration correction approaches [70, 71, 78, 90, 23, 28] correct aberrations through estimating aberration kernels and true images without any prior PSF information. Eboli et al. [23] adopted a small CNN for color correction, and Gong et al. [28] utilized the physical properties of lenses to improve image quality. These blind methods can be applied generally without demanding preprocess (e.g., PSF calibration), but restoration quality may be unsatisfactory due to uncertainties when estimating PSF.
On the one hand, there have been aberration correction models designed for metalens imaging  [80, 9]. Tseng et al. [80] employed CNNs and Wiener filters for restoration, but it struggled to restore the fine details of the image because of the limited receptive field of CNNs. Chakravarthula et al. [9] introduced diffusion models [36, 74] with Wiener filter to manage aberrations, yet it can introduce unrealistic details due to the stochasticity of the diffusion model.
Aberration correction poses unique challenges in the context of metalens imaging, but ViT has not been fully examined to address them. In this paper, we propose leveraging the ViT, adapting attention mechanisms to deal with complex aberration patterns in metalens imaging, to overcome the limitations faced by CNN and diffusion models.
@@ 3 Methods
Fig. 2 depicts the overall architecture of MetaFormer. It consists of Vision Transformer (ViT) as a backbone network for restoration (Sec. 3.2), Multiple Adaptive Filters Guidance (MAFG) (Sec. 3.3), and Spatial and Transposed self-Attention Fusion (STAF) blocks (Sec. 3.4). MetaFormer produces a residual image which is added to the input aberrated image to acquire the aberration-corrected image.
@@ 3.1 Preliminary
Metalens Imaging An imaging formulation using a spatially invariant point spread function (PSF) can be written as follows [^1] While a more general approach involves a spatially varying PSF, this work focuses on restoring aberrated images using a spatially invariant PSF.,
$I_{abr}=I_{c}*PSF+n,$ (1)
where * denotes the convolution operator, and $I_{abr}$ is the observed image, which can be represented as the convolution of a clean image or an object $I_{c}$ with a single PSF uniformly across the entire image. Then the noise $n$, such as shutter and natural noise, is added to generate the captured image $I_{abr}$. Convolution can be replaced with simple element-wise multiplication when $I_{c}$ and $PSF$ are in the frequency domains [8]. The PSF of metalens varies significantly with wavelength because different colors focus at different distances from the metalens. This produces color fringing and a low contrast in the captured image.
3.2 Image Restoration using Vision Transformer
Chromatic aberrations of metalens differ across different wavelengths of light, implying that the aberrations affect each channel of the image distinctively. Vision Transformer (ViT) can model these channel-specific differences, leveraging their multi-head attention to process each channel properly. Additionally, metalens-induced aberrations can degrade the images at various scales, affecting both high-frequency details (e.g., color fringes at sharp edges) and low-frequency components (e.g., gradual intensity shifts). ViT can represent multi-scale features, which can help in correcting different types of distortions.
As discussed, various ViT-based approaches [81, 55, 85, 59, 15, 77, 16, 92] have been developed for restoring images captured by traditional cameras. Among these, Restormer [92] has demonstrated remarkable restoration performance across multiple tasks, utilizing an encoder-decoder architecture with transposed self-attention (TA), a novel feed-forward network with gating mechanism [19], and depth-wise convolutions [37]. Nevertheless, applying Restormer directly to metalens image restoration does not yield satisfactory outcomes, as metalenses introduce more complex aberrations and distortions than conventional lenses, demanding a more sophisticated and tailored solution. Therefore, we employ Restormer as the backbone network and incorporate novel aberration-correcting modules to better align ViT with the specific challenges of metalens imaging. Additionally, we train ViT non-blindly, harnessing point spread function (PSF) for training, as the PSF for metalens can be easily obtained via calibration methods compared to the traditional cameras.
@@ 3.3 Multiple Adaptive Filters Guidance
Before feeding the aberrated input images to the restoration network, we deconvolve the image first with multiple Wiener filters [86]. Wiener filter is a linear filter that has been used to reduce noise in signals, especially in the context of image processing and restoration [20, 21, 96, 98, 69, 35]. It minimizes the mean square error between the estimated and true signal, providing an optimal approach to reconstructing signals degraded by noise. It can also balance noise suppression and detail preservation by exploiting the signal-to-noise ratio (SNR), which makes it particularly effective in image restoration tasks. Given a PSF, the Wiener filter in the frequency domain, $W$, is defined as follows,
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is an indexing operator, and $u$ and $v$ are horizontal and vertical spatial frequency respectively, $H$ is a fourier transformed PSF, $H^{*}$ is a complex conjugation of $H$, and $K$ is an inverse of SNR. Note that each color channel has a different PSF, but we do not explicitly denote the channel for brevity, e.g. $H_{c}$ or $W_{c}$.
@@
Figure 4: The process of Multiple Adaptive Filters Guidance (MAFG). It produces $M$ different representations from the input images using $M$ different Wiener filters.
As shown in  Fig. 3, $K$ plays a pivotal role in penalizing the noise; a greater $K$ leads to a smoother representation, whereas a smaller $K$ can preserve more high-frequency details, albeit with the increased noise. Inspired by this, we propose to use multiple Wiener filters to guide aberration correction with several distinct representations. It is not feasible to obtain an optimal Wiener filter with accurate SNR as noise distribution is unknown in the real world. Instead of estimating the noise distribution, we adopt multiple Wiener filters with different $K$. We use $M$ Wiener filters and deconvolve the input image to yield $M$ different representations—some focused on noise removal, others with fine information. Various representations are fed to the restoration model, and they can enrich the features complementarily, which in turn improves aberration correction.
Furthermore, we extend multiple Wiener filters to Multiple Adaptive Filters Guidance (MAFG), illustrated in Fig. 4, which determines $K$ adaptively considering the image intensity. Image with higher intensity tends to have better signal quality, so brighter channels often experience less noise and are less sensitive to noise [25]. Thus, we penalize noise less and capture more information for bright images by adjusting $K$ with the image intensity. Also, we treat each channel differently to avoid suppressing high-SNR details unnecessarily. This is important when addressing metalens-captured images since metalens exhibit wavelength-dependent aberrations where each color channel suffers differently due to the different chromatic responses. Based on this observation, we slightly modify  Eq. 2 for MAFG as follows:
@@
where $m$ is the index of $M$ different filters ($1\leq m\leq M$), $c_{I}\in\mathbb{R}$ is the average intensity of each color channel, and $k_{m}\in\mathbb{R}$ is a hyperparameter for scaling $c_{I}$. The proposed MAFG is adaptive to the input images, which allows flexible responses to the various images.
@@
where fft and ifft are Fast Fourier Transform (FFT) and the inverse FFT [8], respectively. $\odot$’ is the Hadamard product, $x$ and $y$ are pixel coordinates. $\{I_{dec}^{m}\}_{m=1}^{M}$ are concatenated along the channel dimension and fed to the restoration network. It is also worth mentioning that the MAFG can be applied to any other restoration model, including video restoration, with negligible computational costs.

### v2

# Aberration Correcting Vision Transformers for High-Fidelity Metalens Imaging

@@ Abstract
Metalens is an emerging optical system with an irreplaceable merit in that it can be manufactured in ultra-thin and compact sizes, which shows great promise in various applications. Despite its advantage in miniaturization, its practicality is constrained by spatially varying aberrations and distortions, which significantly degrade the image quality. Several previous arts have attempted to address different types of aberrations, yet most of them are mainly designed for the traditional bulky lens and ineffective to remedy harsh aberrations of the metalens. While there have existed aberration correction methods specifically for metalens, they still fall short of restoration quality. In this work, we propose a novel aberration correction framework for metalens-captured images, harnessing Vision Transformers (ViT) that have the potential to restore metalens images with non-uniform aberrations. Specifically, we devise a Multiple Adaptive Filters Guidance (MAFG), where multiple Wiener filters enrich the degraded input images with various noise-detail balances and a cross-attention module reweights the features considering the different degrees of aberrations. In addition, we introduce a Spatial and Transposed self-Attention Fusion (STAF) module, which aggregates features from spatial self-attention and transposed self-attention modules to further ameliorate aberration correction. We conduct extensive experiments, including correcting aberrated images and videos, and clean 3D reconstruction. The proposed method outperforms the previous arts by a significant margin. We further fabricate a metalens and verify the practicality of our method by restoring the images captured with the manufactured metalens. Code and pre-trained models are available at https://benhenryl.github.io/Metalens-Transformer.
@@ 1 Introduction
Metalens is an advanced optical device that leverages nanoscale engineering to manipulate light. Unlike traditional lenses relying on glass or plastic surfaces, the metalens harnesses a metasurface, a thin optical element consisting of nanostructures modulating optical properties such as phase and amplitude. It can be fabricated in ultra-thin and compact sizes, making it an ideal option for various domains, including the medical industries, consumer electronics, and augmented/virtual reality (AR/VR) [96, 74, 88, 27, 64, 50, 39, 48]. Despite its potential for miniaturization, existing metalenses have faced critical challenges, such as aberrations and distortions in the captured images, hampering their broad uses across diverse applications.
Aberrations in metalens occur due to practical constraints and the physical properties of lens materials. Wavelength-dependent nanostructures of the metasurface make different wavelengths of lights experience different phase shifts, which results in them being focused at different points, leading to chromatic aberration. Furthermore, its spherical surface is susceptible to spherical aberrations where rays transmitting through the edges and center of the lens focus at different locations. Even worse, these aberrations are not uniform, rather varying across the image field, posing a more intricate problem of correcting them.
@@
Several previous works [9, 21, 26] have attempted to correct these aberrations, primarily focusing on conventional camera lenses. While effective in some cases, these approaches fall short in addressing the unique challenges posed by metalens-induced aberrations. More recently, Tseng et al. [80] and Chakravarthula et al. [8] have proposed deep learning-based restoration methods for metalenses. Tseng et al. [80] adopted convolutional neural networks (CNNs) and Chakravarthula et al. [8] employed a diffusion model [32, 73] to correct distortions. While promising, they have yet to reach the desired restoration quality.
Besides the aberration corrections, extensive research has addressed other image degradations such as blur, noise, raindrops, haze, etc. Generative models [34, 62, 42, 43], CNNs [56, 94, 59, 28, 45, 92], or Vision Transformer (ViT) [93, 51, 12, 15, 81] have all been employed for image restoration, demonstrating strong performance. However, directly applying these methods to metalens-induced aberration correction is ineffective due to the aberration patterns of metalens that vary in response to different wavelengths, making them significantly more complex and challenging to correct using standard restoration techniques.
Among various frameworks, ViT-based approaches have a strong potential for metalens image restoration. ViT can effectively capture global context by leveraging the self-attention mechanism, allowing each pixel in an image to communicate with all others. This global receptive field is particularly advantageous for metalens restoration, where understanding the overall structure helps suppress noise and artifacts while preserving fine details. Moreover, self-attention modules enable dynamic weighting of features, encouraging them to learn spatially varying corrections, which is essential in metalens image restoration since the severity of aberration is different across the image. Despite these advantages, ViT-based methods remain largely unexplored for metalens restoration.
In this work, we propose a novel aberration correction Transformer specifically designed for metalens image restoration. First, we suggest a Multiple Adaptive Filters Guidance (MAFG) module, which leverages a set of Wiener filters [86] alongside a cross-attention mechanism to dynamically weight features for more effective aberration correction. MAFG offers two key advantages. First, it is well-suited for handling spatially varying aberrations. MAFG produces features that can consider different degrees of aberration by reweighting multiple features generated with different filters via cross-attention. Second, by employing filters with varying noise suppression parameters, MAFG enhances robustness against unknown aberration distributions, making it adaptable to a wide range of imaging conditions.
Furthermore, we propose a Spatial and Transposed self-Attention Fusion (STAF) module, which integrates both spatial and transposed self-attention to enhance feature aggregation while considering the distinct roles of the encoder and decoder. In STAF, spatial and transposed self-attention operate in parallel, with their features fused to guide the encoder in capturing global context and the decoder in refining fine details. We also encourage attention modules in the decoder to better capture intricate details, further improving restoration quality.
@@
We conducted extensive experiments across various imaging scenarios to validate the effectiveness of the proposed method. On image dataset with spatially varying aberration, it demonstrated notable performance improvements. We further applied the proposed techniques to video restoration and 3D reconstruction tasks to confirm their superiority. Finally, we fabricated metalens using the open-source PSF of a metalens [80] and evaluated our method on the captured images. The corrected images demonstrated promising results, revealing new possibilities in metalens imaging. To sum up, our contributions are as follows:
  • We propose a novel neural architecture for metalens image restoration, achieving state-of-the-art performance with a 6.45 dB improvement in PSNR over the baseline model.
@@ 2 Related Works
Image Restoration Models
Image restoration aims to recover high-quality images from degraded observations. Traditionally, Wiener deconvolution [86] has been used to recover images by solving inverse problems. However, it often struggled with noise and complex degradations, limiting their real-world applications. Recently, deep learning-based methods have shown great success in image restoration tasks in a data-driven manner. Vision Transformer (ViT) [20], in particular, has achieved remarkable performances in various image restoration tasks, including denoising [81, 10, 85, 24, 93, 49], deblurring [85, 93, 55, 41, 79], and super resolution [13, 14, 51, 77, 57, 15], by exploiting self-attention, capturing global interactions effectively.
Restormer [93] stood out with its superior performance, introducing transposed self-attention mechanisms that apply self-attention across the channel of an image, enabling efficient processing of high-resolution images. While this approach enhances computational efficiency, it may limit the model’s ability to fully exploit spatial dependencies. To address this, X-Restormer [12] incorporated spatial self-attention mechanisms allowing for better capture of spatial relationships. In this work, building upon these advancements, we propose a novel architecture that further enhances restoration quality for metalens images.
Figure 2: The overview of our method. It comprises Multiple Adaptive Filters Guidance (MAFG) which produces different representations with various noise-detail balances, and a Spatial and Transposed self-Attention Fusion (STAF) module that aggregates features differently in encoder and decoder.
Aberration Correction Models
@@
Blind aberration correction approaches [68, 69, 78, 91, 21, 26] correct aberrations by estimating aberration kernels without any prior PSF information. Eboli et al. [21] adopted a small CNN for color correction, and Gong et al. [26] utilized the physical properties of lenses to improve image quality. These blind methods can be applied generally without demanding preprocess (e.g., PSF calibration), but restoration quality may suffer due to uncertainties in accurately estimating the PSF.
Several aberration correction models have been specifically designed for metalens imaging [80, 8]. Tseng et al. [80] employed CNNs and Wiener filters for restoration but faced challenges in preserving fine image details. This limitation may arise from the restricted receptive field of CNNs. Chakravarthula et al. [8] introduced diffusion models [32, 73] with Wiener filters to manage aberrations, yet it can generate unrealistic details due to the stochasticity of the diffusion, compromising the overall image quality.
Aberration correction presents unique challenges in metalens imaging, yet the potential of Vision Transformers (ViT) has not been fully explored for this task. In this paper, we propose a novel approach that leverages ViT, developing new architectural modules and adapting its attention mechanisms to effectively handle the complex aberration patterns in metalens imaging.
@@ 3 Methods
Fig. 2 depicts the overall architecture of the proposed method. It consists of Vision Transformer (ViT) as a backbone network for restoration (Sec. 3.2), Multiple Adaptive Filters Guidance (MAFG) (Sec. 3.3), and Spatial and Transposed self-Attention Fusion (STAF) blocks (Sec. 3.4).
@@ 3.1 Preliminary
Metalens Imaging An imaging formulation using a spatially varying point spread function (PSF) can be written in patch-level imaging as follows,
$I_{abr}[h,w]=I_{c}[h,w]*PSF[h,w]+n[h,w],$ (1)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is an indexing operator, * denotes the convolution operator, $h\in\{1,2,...,H\}$ and $w\in\{1,2,...,W\}$ are the indices of $H\times W$ image and PSF patches. $I_{abr}[h,w]$ is the observed image, which can be represented as the convolution of an object $I_{c}[h,w]$ with a corresponding PSF at patch $[h,w]$. Then the noise $n[h,w]$, such as shutter and natural noise, is added to generate the captured image $I_{abr}$. The PSF of metalens varies significantly with wavelength because different colors focus at different distances from the metalens. This produces color fringing and a low contrast in the captured image. Such optical degradations vary gradually across the field of view, implying the distortions are similar within a small neighborhood; hence, the imaging can be formulated at a patch level.
3.2 Vision Transformer for Metalens Images
As discussed, various ViT-based approaches [81, 51, 85, 55, 14, 77, 15, 93] have been developed for restoring images captured by traditional cameras. While ViT has not been fully explored for metalens image restoration, it inherently has promising capabilities to handle the severe aberrations present in metalens images. First, ViT can adaptively adjust its attention weight to focus on different areas of the image, making it particularly suited for handling the spatially varying aberrations of metalens imaging. Furthermore, chromatic aberrations of metalens differ across different wavelengths of light, causing each image channel to be affected differently. ViT’s multi-head attention mechanism allows it to model these channel-specific discrepancies, processing each channel effectively. Lastly, metalens-induced aberrations can degrade the images at various scales, affecting both high-frequency details (e.g., color fringes at sharp edges) and low-frequency components (e.g., gradual intensity shifts). ViT’s ability to capture multi-scale features enhances its capacity to correct different types of distortions.
Such capabilities motivated us to adopt ViT for metalens image restoration. Nevertheless, applying ViT directly to metalens image restoration is ineffective, as metalenses introduce more complex aberrations and distortions than conventional lenses, demanding a more specialized solution. Therefore, we employ ViT as the backbone network and incorporate novel aberration-correcting modules to better align ViT with the specific challenges of metalens imaging.
@@ 3.3 Multiple Adaptive Filters Guidance
Before feeding the aberrated images to the restoration network, we deconvolve the image first with multiple Wiener filters [86]. Wiener filter is a linear filter that has been used to reduce noise in signals, in the context of image processing and restoration [18, 19, 97, 99, 67, 31]. It can balance noise suppression and detail preservation by exploiting the signal-to-noise ratio (SNR), which makes it particularly effective in image restoration tasks. Given a PSF, the Wiener filter in the frequency domain, $W$, is defined as follows,
@@
where $u$ and $v$ are horizontal and vertical spatial frequency respectively, $H$ is a fourier transformed PSF, $H^{*}$ is a complex conjugation of $H$, and $K$ is an inverse of SNR. Note that each color channel has a different PSF, but we do not explicitly denote the channel for brevity, e.g. $H_{c}$ or $W_{c}$.
@@
Figure 4: The process of Multiple Adaptive Filters Guidance (MAFG). It produces $M$ different representations from the input images using $M$ different Wiener filters and applies cross-attention to reweight them.
As shown in  Fig. 3, $K$ plays a pivotal role in penalizing the noise; a greater $K$ leads to a smoother representation, whereas a smaller $K$ can preserve more high-frequency details, albeit with the increased noise. Inspired by this, we propose to use multiple Wiener filters to guide aberration correction with several distinct representations. It is not feasible to obtain an optimal Wiener filter with accurate SNR as noise distribution is unknown in the real world. Such design can approximate the ideal filter by employing $M$ filters with different $K$ in a wide range.
Furthermore, we extend multiple Wiener filters to Multiple Adaptive Filters Guidance (MAFG), illustrated in Fig. 4, which determines $K$ adaptively considering the image intensity. Image with higher intensity tends to have better signal quality, so brighter channels often experience less noise [23]. Thus, we penalize noise less and capture more information for bright images by adjusting $K$ with the image intensity. Also, we treat each channel differently to avoid suppressing high-SNR details unnecessarily, since metalens exhibit wavelength-dependent aberrations where each color channel suffers differently due to the different chromatic responses. Based on this observation, we slightly modify  Eq. 2 for MAFG as follows:
@@
where $m\in\{1,\dots,M\}$ is the index of $M$ different filters, $c_{I}\in\mathbb{R}$ is the average intensity of each color channel, and $k_{m}\in\mathbb{R}$ is a hyperparameter for scaling $c_{I}$. The proposed MAFG is adaptive to the input images, which allows flexible responses to the various images.
@@
where fft and ifft are Fast Fourier Transform (FFT) and the inverse FFT [7], respectively. $\odot$’ is the Hadamard product, $x$ and $y$ are pixel coordinates.
Then, they are processed by a cross-attention block, where the query is $I^{k_{med}}_{dec}$, the deconvolved image with a median of $k_{m}$, which is a compromise between noise suppression and detail preservation. Cross-attention can reweight the features, which it can attend more to the filter outputs with high $k_{m}$ (severe aberration, red boxes in Fig. 3) or it can favor the filter outputs with low $k_{m}$ (mildly aberrated parts, blue boxes in Fig. 3), resulting in spatially adaptive correction. This is especially important for metalens images, where aberration occurs non-uniformly. The reweighted feature is then fed to the restoration network.